# Foundation Model of Electronic Medical Records for Adaptive Risk Estimation

### v3

@@ Acknowledgments
LR 0.677 [0.661, 0.693] 0.160 [0.138, 0.179] 0.571 [0.535, 0.645] 0.683 [0.626, 0.712]
RF 0.672 [0.654, 0.680] 0.147 [0.125, 0.158] 0.570 [0.552, 0.652] 0.688 [0.599, 0.690]
GB 0.698 [0.681, 0.713] 0.165 [0.147, 0.181] 0.616 [0.576, 0.673] 0.667 [0.609, 0.709]
MLP 0.694 [0.681, 0.706] 0.162 [0.144, 0.180] 0.628 [0.591, 0.659] 0.660 [0.615, 0.690]
Med2Vec 0.656 [0.646, 0.670] 0.133 [0.121, 0.150] 0.564 [0.547, 0.600] 0.663 [0.624, 0.679]
LSTM 0.689 [0.676, 0.702] 0.168 [0.147, 0.183] 0.654 [0.574, 0.661] 0.612 [0.609, 0.679]
AutoScore 0.624 [0.612, 0.636] 0.074 [0.068, 0.082] 0.589 [0.548, 0.635] 0.601 [0.563, 0.640]
MEDS-Tab 0.696 [0.679, 0.713] 0.163 [0.143, 0.185] 0.635 [0.598, 0.680] 0.661 [0.633, 0.704]
ETHOS (ours) 0.740 [0.723, 0.760] 0.199 [0.171, 0.230] 0.659 [0.643, 0.678] 0.696 [0.685, 0.708]
@@
      Longest 221,122 106,936 221,122
      Q3 1,041 1,052 1,042
      Median 322 331 323
      Mean 1,249 1,259 1,250
      Q1 114 115 114
@@
    diagnosis icd_version, icd_code, stay_id
    edstays intime, arrival_transport, hadm_id stay_id, outtime, disposition hadm_id, stay_id
    pyxis name, charttime, stay_id
    triage acuity, stay_id
    vitalsign temperature, charttime, stay_id heartrate, charttime, stay_id resprate, charttime, stay_id o2sat, charttime, stay_id sbp, dbp, charttime stay_id, pain, charttime stay_id
@@
    admissions admission_type, admission_location, admittime insurance, marital_status, race hadm_id, discharge_location, dischtime hadm_id
    diagnoses_icd icd_version, icd_code, hadm_id
    drgcodes drg_type, drg_code, description hadm_id
    emar medication, event_txt, charttime hadm_id, emar_id, emar_seq
    hcpcsevents short_description, hadm_id, chartdate
    labevents itemid, valueuom, hadm_id charttime, valuenum, value
    omr result_name, result_value, chartdate
    patients gender, dod
    procedures_icd icd_version, icd_code, hadm_id chartdate
    transfers eventtype, careunit, intime hadm_id
@@
    icustays first_careunit, intime, hadm_id stay_id, last_careunit, outtime hadm_id, stay_id
@@ A. Monte Carlo Estimator
    Define an indicator function $I(\mathbf{x}^{(i)}\in A)$, which is $1$ if the sample $\mathbf{x}^{(i)}$ lies in $A$, and 00 otherwise.

### v4

@@ Acknowledgments
LR 0.679 [0.661, 0.697] 0.161 [0.141, 0.185] 0.562 [0.544, 0.645] 0.699 [0.613, 0.718]
Med2Vec 0.621 [0.601, 0.640] 0.128 [0.110, 0.148] 0.560 [0.477, 0.595] 0.615 [0.568, 0.725]
RF 0.673 [0.655, 0.691] 0.150 [0.131, 0.173] 0.642 [0.549, 0.665] 0.599 [0.594, 0.693]
GB 0.697 [0.680, 0.714] 0.165 [0.143, 0.188] 0.623 [0.592, 0.704] 0.662 [0.582, 0.690]
MLP 0.693 [0.675, 0.711] 0.168 [0.147, 0.192] 0.603 [0.579, 0.676] 0.675 [0.607, 0.701]
LSTM 0.689 [0.671, 0.708] 0.164 [0.143, 0.186] 0.595 [0.566, 0.653] 0.680 [0.633, 0.722]
MEDS-Tab 0.714 [0.696, 0.731] 0.189 [0.167, 0.214] 0.645 [0.575, 0.689] 0.657 [0.617, 0.742]
ETHOS (ours) 0.745 [0.728, 0.762] 0.214 [0.190, 0.239] 0.669 [0.611, 0.698] 0.685 [0.657, 0.757]
@@
    Longest 221,346 107,147 221,346
    Q3 1,053 1,063 1,055
    Median 331 340 332
    Mean 1,262 1,272 1,263
    Q1 122 123 122
@@
    diagnosis
icd_code, icd_version, stay_id
    edstays
arrival_transport, disposition, hadm_id
intime, outtime, stay_id
    pyxis
charttime, name, stay_id
    triage
acuity, dbp, heartrate
o2sat, pain, resprate
sbp, stay_id, temperature
    vitalsign
charttime, dbp, heartrate
o2sat, pain, resprate
sbp, stay_id, temperature
@@
    admissions
admission_location, admission_type, admittime
discharge_location, dischtime, hadm_id
insurance, marital_status, race
    diagnoses_icd
hadm_id, icd_code, icd_version
    drgcodes
description, drg_code, drg_type
hadm_id
    emar
charttime, emar_id, event_txt
hadm_id, medication
    hcpcsevents
chartdate, hadm_id, short_description
    labevents
charttime, hadm_id, itemid
valuenum, valueuom
    omr
chartdate, result_name, result_value
    patients
dod, gender
    procedures_icd
chartdate, hadm_id, icd_code
icd_version
    transfers
careunit, eventtype, hadm_id
intime
@@
    icustays
first_careunit, hadm_id, intime
last_careunit, outtime, stay_id
Table S8: Side-by-side view of selected columns from the original sample tables (sourced from MIMIC-IV-DEMO) compared with the format of the tokenized timelines in ETHOS.
Original data Tokenized data
patients.csv timeline.csv
subject_id gender anchor_age subject_id time code
10038081 F 63 10000248 2192-11-29 18:44:00 ED_REGISTRATION
10019917 M 44 10000248 2192-11-29 18:44:00 ED_TRANSPORT//WALK_IN
10019568 F 59 10000248 2192-11-29 18:44:00 ED_ACUITY
10031404 F 82 10000248 2192-11-29 18:44:00 Q2
10008287 F 43 10000248 2192-11-29 19:03:00 15m-45m
admissions.csv 10000248 2192-11-29 19:03:00 LAB//51146//%
subject_id hadm_id admittime 10000248 2192-11-29 19:03:00 Q3
10035631 22732862 2112-11-10 15:55:00 10000248 2192-11-29 19:03:00 LAB//51200//%
10020786 23488445 2189-06-09 12:45:00 10000248 2192-11-29 19:03:00 Q7
10020187 26842957 2170-02-24 00:00:00 10000248 2192-11-29 19:03:00 LAB//51221//%
10005866 27167814 2148-03-10 16:16:00 10000248 2192-11-29 19:03:00 Q4
10002428 28676446 2157-07-16 04:09:00 10000248 2192-11-29 19:03:00 LAB//51222//G/DL
diagnoses.csv 10000248 2192-11-29 19:03:00 Q5
subject_id hadm_id icd_code 10000248 2192-11-29 19:03:00 LAB//51144//%
10016742 29281842 N179 10000248 2192-11-29 19:04:00 Q6
10023117 21607814 I428 10000248 2192-11-29 19:37:00 15m-45m
10040025 25933959 I130 10000248 2192-11-29 19:37:00 HOSPITAL_ADMISSION
10014354 27494880 Z955 10000248 2192-11-29 19:37:00 OBSERVATION
10014354 27487226 I5033 10000248 2192-11-29 19:37:00 INSURANCE_MEDICAID
Table S9: Estimated energy consumption of training ETHOS compared to training large language models. ETHOS is a dedicated model specifically designed for the electronic health records (EHR) domain, which allows it to be substantially smaller and more efficient to train than general-purpose large language models. With only 45 million parameters, ETHOS was trained on 8 A100 GPUs over 46 hours, consuming an estimated 220 kWh of energy. In contrast, universal LLMs such as GPT-3 (175B parameters), LLaMA 3 (8B), and Falcon (40B) require orders of magnitude more compute and energy, consuming between 307,000 and over 1.2 million kWh.
Model Name Params (B) GPUs Used Duration Est. GPU-hours Est. Energy (kWh)
ETHOS 0.045 8×A100 46 hours 368 $\sim$220
GPT-3 175 1024×V100 $\sim$34 days $\sim$835,000 $\sim$1,287,000
LLaMA 3 (8B) 8 $\sim$512×A100 $\sim$25 days $\sim$307,000 $\sim$490,000
Falcon (40B) 40 384×A100 $\sim$21 days $\sim$193,000 $\sim$307,000
@@
Figure S3: AUPRC Curves for ETHOS Across All Prediction Tasks. Precision-recall (PR) curves and corresponding area under the precision-recall curve (AUPRC) values are shown for seven prediction tasks: Hospital Mortality, ICU Admission, Prolonged Stay (>10 days), Composite Outcome (Hospital Mortality + ICU Admission + Prolonged Stay), Hospitalization at Triage, Critical Outcome Within 12h at Triage, and Emergency Department (ED) Re-presentation Within 72h. Each plot includes the PR curve (orange) and unique thresholds (crosses). ETHOS shows good precision-recall performance across several tasks, with AUPRC values ranging from 0.199 (ED Re-presentation) to 0.887 (Hospitalization at Triage), reflecting the class imbalance present in each task.
@@ A. Monte Carlo Estimator
    Define an indicator function $I(\mathbf{x}^{(i)}\in A)$, which is $1$ if the sample $\mathbf{x}^{(i)}$ lies in $A$, and 0 otherwise.
@@ C. Convergence by the Law of Large Numbers
Appendix B Intuitive Operation of ETHOS
Figure S1 shows the decoder-only transformer backbone of ETHOS, which follows the standard GPT design. Below we provide an intuitive, step-by-step explanation of how ETHOS works “zero-shot,” i.e. with no additional fine-tuning for specific prediction tasks.
1. Treating Patient History as a “Story.”
Each Patient Health Timeline (PHT) is a long sequence of discrete tokens, think of each token as a “word” in a clinical narrative. Static tokens (e.g., demographics) set the scene, categorical tokens (e.g., diagnoses) are like domain-specific vocabulary, and time-interval tokens keep track of elapsed time between events. Just as a language model reads a sentence and predicts the next word, ETHOS reads a patient’s history and predicts the next medical event token.
2. Zero-Shot Generative Inference.
Because ETHOS has been trained to model the joint distribution of these tokens, at inference time it can continue any PHT without further training. We provide the model with the tokens of a patient’s past, and then let it generate new tokens one at a time. No task-specific labels or fine-tuning are needed—hence “zero-shot.” Each new token corresponds to a plausible future event (e.g., a lab draw, medication change, or hospital admission).
3. Exploring Possible Futures via Sampling.
To capture uncertainty, we perform many independent “completions” of the PHT. Each completion (an fPHT) is analogous to asking, “What could happen next?” and letting the model write the next chapter of the patient’s story. We use nucleus sampling (top-$p$) to allow diversity: at each step, ETHOS randomly selects the next token from the smallest set whose cumulative probability exceeds $p=0.9$. By generating $N$ such fPHTs, we build a Monte Carlo ensemble of future trajectories.
4. From Generated Tokens to Risk Estimates.
Some tokens are can be treated as outcome (e.g., mortality). To compute inpatient mortality, we simply count how many of the $N$ simulated fPHTs include that token while patient is in hospital (before discharge token). The fraction $M/N$ then provides a direct, probabilistic estimate of risk. This simulation-based approach naturally accounts for multiple pathways and branching possibilities in a patient’s course.
5. Toward Controlled Cohort Generation.
Because ETHOS is a pure generator, one can modify the sampling distribution, reweighting certain token types (e.g., age or comorbidity tokens), to synthesize patient cohorts with desired characteristics. This capability opens the door to fairness-aware risk modeling and targeted “what-if” analyses, which we plan to explore in future work.
Appendix C Baseline Models
To demonstrate the applicability of ETHOS as the base for ARES we compared its performance on three benchmark tasks against other baseline models. Those tasks were: prediction of the hospital admission at triage, prediction of the critical outcome (death or transfer to ICU within 12 hours) at triage, and ED re-presentation within 72 hours after discharge from ED. We followed the baseline models utilized for the benchmark tasks as presented in [29]. Firstly, we used clinically applied scoring systems: Modified Early Warning Score (MEWS) [13], National Early Warning Score (NEWS, versions 1 and 2) [12, 30, 31], Rapid Emergency Medicine Score (REMS) [32], and Cardiac Arrest Risk Triage (CART) [33]. Those scores require the collection of specific clinical features like Heart Rate, Respiratory Rate or Oxygen Saturation. They cannot adapt to the dynamic state of the patient unless all the features used by the scoring system are remeasured. Similarly, Emergency Severity Index (ESI) [34], a five-level triage system, was assigned by a nurse. In addition, we created custom scores with the AutoScore method that generates scoring systems from clinical features automatically in six steps, including variable ranking and transformation as well as score derivation by weighting and normalization  [43]. Next, we utilized classic machine learning algorithms, including Logistic Regression (LR), which is a linear classifier, Multi-layer Perceptron (MLP), a non-linear classifier and tree-based, ensemble algorithms Random Forest (RF) and Gradient Boosting (GB). Finally, we applied more advanced deep learning-based algorithms Med2Vec [35], which applies non-linear transformations to vector embeddings of ICD codes and Long Short-Term Memory (LSTM) [36] a neural network accounting for temporal changes in data for the ED re-presentation task. We make our adapted baseline model training code publicly available for reproducibility purposes  [^2] https://github.com/ipolharvard/mimic4ed-benchmark.
Details of model training
ETHOS was implemented as a decoder-only transformer. Training data were constructed by appending an “End of timeline” token to each patient’s PHT and concatenating all timelines into one long token sequence. The model was then trained, in an unsupervised fashion, to predict the next token given its preceding context by minimizing the standard cross-entropy loss, exactly as in large-scale language-model pretraining. Given the scale of our PHT corpus (hundreds of millions of tokens) and the model’s complexity, we matched the parameter count of GPT-2 as a starting point, conducted a focused hyperparameter search (see Figure S1), and made heuristic adjustments to optimize convergence. Training was performed across eight NVIDIA GPUs over approximately 36 hours, reflecting resource requirements similar to those for open-domain transformer models. AdamW optimizer (decoupled weight decay) with hyperparameters matched to those in nanoGPT/GPT-2–scale pretraining. In our experiments, we used a constant learning rate of $4e^{-4}$, weight-decay of 0.01, and default momentum parameters ($\beta_{1}=0.9$ and $\beta_{2}=0.999$ training for 100K steps per federated round with batch size 32. For more details, refer to [15]. All code including training scripts is publicly available in our GitHub repository at https://github.com/ipolharvard/ethos-ares.